# Incremental Sequence Labeling: A Tale of Two Shifts

### v1

@@ Abstract
The incremental sequence labeling task involves continuously learning new classes over time while retaining knowledge of the previous ones. Our investigation identifies two significant semantic shifts: E2O (where the model mislabels an old entity as a non-entity) and O2E (where the model labels a non-entity or old entity as a new entity). Previous research has predominantly focused on addressing the E2O problem, neglecting the O2E issue. This negligence results in a model bias towards classifying new data samples as belonging to the new class during the learning process. To address these challenges, we propose a novel framework, Incremental Sequential Labeling without Semantic Shifts (IS3). Motivated by the identified semantic shifts (E2O and O2E), IS3 aims to mitigate catastrophic forgetting in models. As for the E2O problem, we use knowledge distillation to maintain the model’s discriminative ability for old entities. Simultaneously, to tackle the O2E problem, we alleviate the model’s bias towards new entities through debiased loss and optimization levels. Our experimental evaluation, conducted on three datasets with various incremental settings, demonstrates the superior performance of IS3 compared to the previous state-of-the-art method by a significant margin.
@@ 1 Introduction
The conventional sequence labeling task typically involves categorizing data into a predetermined set of fixed categories Lample et al. (2016). However, this approach may need to be revised in natural language processing scenarios, such as the named entity recognition task, where new types of entities continuously emerge. Adapting a fixed set of categories becomes challenging when faced with the dynamic nature of new entity classification requirements. Consequently, continuous model updates are essential to accommodating evolving entity types. Previous studies have advocated for adopting continual learning Parisi et al. (2019); Monaikul et al. (2021), also known as lifelong learning or incremental learning. Continual learning is a paradigm designed to train models capable of adapting to the continual addition of new categories in real-world scenarios while ensuring that knowledge of old categories is retained. For instance, voice assistants like Siri frequently encounter new event types, such as pandemics, to better understand users’ latest intentions and provide information on health protection Monaikul et al. (2021).
@@
Due to constraints imposed by storage limitations and privacy concerns, there exists a shortage of training data about the older categories He and Zhu (2022). Additionally, the manual relabeling of all categories within the new training dataset would incur substantial costs and time investment De Lange et al. (2021); Bang et al. (2021). Consequently, the model undergoes continuous updates using a freshly acquired dataset comprising the new categories. As depicted in Fig.1, the model undergoes training based on the current ground-truth label and undergoes testing using the full ground-truth label.
The incremental sequence labeling task faces a significant challenge known as the catastrophic forgetting problem, as extensively discussed in previous studies McCloskey and Cohen (1989); Robins (1995); Goodfellow et al. (2013); Kirkpatrick et al. (2017). This issue manifests as semantic shifts, leading to a decrease in the discriminative power of entity classes Zhang et al. (2023b); Ma et al. (2023). In this paper, we decompose the problem into two primary semantic shifts in the incremental sequence task: E2O and O2E. The first semantic shift, E2O, arises from the presence of non-entities, potential old entities (mislabelled as non-entities), and new entities in the new dataset. Progress has been made in addressing E2O through methods falling into three categories: (1) Methods based on knowledge distillation: For instance, RDP proposes a knowledge distillation loss incorporating inter-task relations Zhang et al. (2023b). At the same time, CFPD introduces a pooled feature distillation loss to alleviate catastrophic forgetting Zhang et al. (2023a). (2) Methods based on pseudo-labels: OCILNER utilizes class prototypes to label new data Ma et al. (2023), and CPFD employs old models to label predictions of new data. (3) Methods based on freezing models: Examples include ICE Liu and Huang (2023), which freezes the backbone model and old classifiers to maintain the stability of the old classes at the expense of learning new classes.
@@ 2 Related Work
Incremental Learning The model continually acquires new tasks intending to achieve optimal performance on tasks previously learned Gepperth and Hammer (2016); Wu et al. (2019); van de Ven et al. (2022). There are three main categories of current incremental learning methods: regularization-based, rehearsal-based, and architecture-based. Regularization-based methods place constraints on model weights Kirkpatrick et al. (2017); Zenke et al. (2017), representations of intermediate layer features Hou et al. (2019); Douillard et al. (2020), and output probabilities Li and Hoiem (2017). Rehearsal-based methods overcome forgetting by saving some of the data containing the old classes for learning with the new classes Lopez-Paz and Ranzato (2017); Shin et al. (2017). Alternatively, architecture-based approaches involve dynamically expanding the network structure to allow for more data as new classes are added Hou et al. (2018); Yan et al. (2021).
@@
Methods for incremental sequence labeling tasks can be categorized into distillation-based, rehearsal-based, and other approaches. Distillation-based methods encompass ExtendNER Monaikul et al. (2021), which is the pioneer in applying knowledge distillation to incremental sequence labeling task, RDP Zhang et al. (2023b) with a relational distillation approach, and CPFD Zhang et al. (2023a) utilizing pooled features distillation loss. CFNER Zheng et al. (2022) introduces a causal framework for extracting new causal effects in entities and non-entities. Rehearsal-based approaches include KCN Cao et al. (2020) and KD+R+K Yu et al. (2021), both employing rehearsal samples to address class imbalance and catastrophic forgetting in incremental event detection. L&R Xia et al. (2022) proposes a learn-and-review framework by training a new backbone model and a generative model simultaneously, generating synthetic samples of the old class to be trained with new samples. OCILNER Ma et al. (2023) uses rehearsal samples to compute class feature centers as class prototypes, generates an entity-oriented feature space through comparative learning, and annotates new data with pseudo-labels using class prototypes. Other methods encompass span-based and freezing model-based approaches, among others.
@@ 3 Problem Formulation
Formally, the objective of incremental sequence labeling is to acquire knowledge through a series of tasks $\mathcal{T}=\{\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{N}\}$. Each task contains its own dataset $\mathcal{D}_{t}=\{(x^{i},y^{i})|y^{i}\in\mathcal{Y}_{t}\}$ where $(x^{i},y^{i})$ is a pair formed by the input token sentence and the label corresponding to each token in the sentence and $\mathcal{Y}_{t}$ stands for the current label set. Notably, $y^{i}$ only labels the token corresponding to the current task $t$, and the other tokens are labeled as O class (potential old entities $\mathcal{Y}_{1:t}$, and unseen entities $\mathcal{Y}_{t+1:N}$). At task $t$ ($t>1$), the new model $\mathcal{M}_{t}$ learns only from the new dataset and is expected to perform well on the learned classes $\bigcup_{i=1}^{t}\mathcal{Y}_{i}$.
@@ 4 Method
In this section, we systematically address the catastrophic forgetting problem by decomposing it into two distinct semantic shift challenges (Section 4.1). Subsequently, we present a comprehensive framework designed to address these semantic shifts individually, focusing on E2O in Section 4.2 and O2E in Section 4.3. The overarching goal is to effectively mitigate the catastrophic forgetting problem, as illustrated in Fig.4.
@@ 4.1 Two semantic shift problems
E2O refers to the model incorrectly categorizing entities as non-entities during the learning process. This misclassification stems from the incremental sequence labeling task, where only new entities are labeled in the new dataset, potentially causing old entities to be erroneously labeled as non-entities. For instance, in Fig.3, the name "Amy" is mistakenly labeled as a non-entity. This misclassification induces a gradual shift in the semantics of old entities towards non-entities, leading to a blurred boundary between the two classes. Several previous approaches have addressed this bias issue. Methods like RDP focus on designing improved distillation techniques to maintain the stability of the model’s old entities. Similarly, OCILNER utilizes comparative learning to obtain a more discriminative feature space, clarifying the classification boundaries between entities and non-entities. These strategies aim to mitigate the impact of E2O, ensuring a more accurate preservation of entity semantics during incremental sequence labeling tasks.
@@
O2E signifies the model incorrectly labeling non-entities or old entities as new entities during the learning process. As seen in Fig.3, our observations indicate that while the model maintains good discrimination between old entities. However, in Fig.2, there is a bias towards new entities in predictions during incremental learning. Our research identifies two key contributing factors to this bias.
@@ 4.2 Solving E2O problem via knowledge distillation
Through Eq.1, "Amy" in Fig.4 corrects the current model’s incorrect labeling via the output probabilities provided by the old model, thus maintaining discriminative properties between old entities.
@@ 4.3.1 Debiasing in Ordinary Cross Entropy
where $y_{i}$ denotes the label of the new entity for the current incremental step $t$. Fig.2 shows that confusion matrix of previous method at incremental step 4. It clearly shows that most predictions are biased towards the recent entity (class 4). We find that such a bias can be found in the cross-entropy loss function. When learning new entities, the model’s gradient update for old entities is defined as:
@@ 4.3.2 Learning with Prototypes
The class prototypes of old entities serve two essential purposes: firstly, they participate in optimization alongside new entities in each batch, ensuring a balanced optimization process among entities. Secondly, these class prototypes act as anchors in the feature space, mitigating the issue of over-labeling new entities. As depicted in Fig.4, the introduction of old prototypes reduces the potential over-labeling of new entities, enhancing the precision of new entity learning.
@@
Our approach differs from OCILNER’s approach which uses prototypes in the following two ways:(1) OCILNER’s approach stores old samples for calculating prototypes. Yet in this paper, we only use the training data in each incremental step for calculating prototypes, and do not introduce replay samples. (2) OCILNER uses prototypes to label new datasets and adopts a cosine similarity as the threshold for entity labeling. However, in this paper, we found that some of the real non-entities also have a high cosine similarity with entities, which can easily produce wrong labeling for real non-entities and exacerbate semantic migration from entities to non-entities.
@@
Table 1: Comparisons with state-of-the-art methods on i2b2. The best results are highlighted in bold and the second best results are underlined. The average of each incremental step is provided in Fig.10.
@@ 5.1 Experimental Setup
Settings We sort the above slices according to initial letter and train them in a FG-a-PG-b manner. FG means that the pre-trained model is trained with a entity types as the initial model and PG means that the initial model is trained with b entity types at each following incremental step.
@@
Figure 5: The comparison between our method and previous state-of-the-art methods on nine incremental learning settings. We report the MacroF1 score after learning the final task. The detailed results are provided in Table 1, Table 7 and Table 5.
Implementation Details We use bert-base-cased model from HuggingFace Wolf et al. (2019) as backbone, with a hidden dimension of $d=768$. We use the AdamW Loshchilov and Hutter (2018) optimizer, with learning rate $1e^{-6}$ and $1e^{-3}$ for backbone and classifier. We report the mean and standard deviation results over five runs.
@@ 5.2 Results and Analysis
Comparisons with State-Of-The-Art To validate the effectiveness of our approach, we conducted exhaustive experiments on the i2b2, OntoNotes5, and MAVEN datasets. We used the Finetune Only (FT) approach as a lower bound for comparison. Table 1 displays the results of the experiments conducted on i2b2. Due to space limitations, we provide the results on MAVEN in Table 5 and OntoNotes5 in Table 7. We show the experimental results under nine incremental learning settings in detail through Fig 5. Our method consistently outperforms the previous state-of-the-art method in multiple settings, from FG-1-PG-1 to FG-8-PG-2. The poor performance of the previous method may be attributed to the ignorance of O2E.
@@
As shown in Fig.6, during the learning process, the previous method, ExtendNER, confuses new entities with non-entities due to O2E and old entities with non-entities due to E2O. Both of them together lead to poor prediction results of the model. We have effectively mitigated the above problems through our framework IS3, which strikes a good balance between maintaining old entities and learning new ones.
To further demonstrate the effectiveness of our method, we visualize the feature representation through T-SNE Van der Maaten and Hinton (2008). As shown in Fig.11 in appendix E, the ExtendNER method faces serious E2O and O2E problems, with new entities and non-entities overwriting old ones, leading to catastrophic forgetting. Our method successfully addresses the issue of semantic bias that arises when the model learns a new task.
@@
Hyper-Parameter Analysis Fig.8 shows the results of different hyper-parameter choices on OntoNotes5 with the setting FG-1-PG-1. We consider two hyper-parameters: the correction factor in the debiased cross-entropy loss $\delta$ and the weight of the prototype loss $\beta$. The results show that $\delta$ around 0.5 reaches the best result, indicating that a moderate penalty effect reduction favors model performance. As $\beta$ keeps increasing, it makes the model overfit for the old prototype, leading to a decrease in model performance.
Case Study We provide an example in Fig.7 to demonstrate that the previous method suffers from an O2E offset when learning a new entity ORG, overwriting the old entity EVENT as a new entity. Simultaneously, the model inherits past O2E issues (labeling [O] as [DATE]). Additionally, it suffers from E2O, which fails to recognize the old entity accurately. Our method effectively balances these two types of offset problems and is more conducive to model learning.

### v2

@@ Abstract
The incremental sequence labeling task involves continuously learning new classes over time while retaining knowledge of the previous ones. Our investigation identifies two significant semantic shifts: E2O (where the model mislabels an old entity as a non-entity) and O2E (where the model labels a non-entity or old entity as a new entity). Previous research has predominantly focused on addressing the E2O problem, neglecting the O2E issue. This negligence results in a model bias towards classifying new data samples as belonging to the new class during the learning process. To address these challenges, we propose a novel framework, Incremental Sequential Labeling without Semantic Shifts (IS3). Motivated by the identified semantic shifts (E2O and O2E), IS3 aims to mitigate catastrophic forgetting in models. As for the E2O problem, we use knowledge distillation to maintain the model’s discriminative ability for old entities. Simultaneously, to tackle the O2E problem, we alleviate the model’s bias toward new entities through debiased loss and optimization levels. Our experimental evaluation, conducted on three datasets with various incremental settings, demonstrates the superior performance of IS3 compared to the previous state-of-the-art method by a significant margin. The data, code, and scripts are publicly available  [^1] https://github.com/zzz47zzz/codebase-for-incremental-learning-with-llm.
@@ 1 Introduction
The conventional sequence labeling task typically involves categorizing data into a predetermined set of fixed categories Lample et al. (2016). However, this approach may need to be revised in natural language processing scenarios, such as the named entity recognition task, where new types of entities continuously emerge. Adapting a fixed set of categories becomes challenging when faced with the dynamic nature of new entity classification requirements. Consequently, continuous model updates are essential to accommodating evolving entity types. Previous studies have advocated for adopting continual learning Parisi et al. (2019); Monaikul et al. (2021), also known as lifelong learning or incremental learning. Continual learning is a paradigm designed to train models capable of adapting to the continual addition of new categories in real-world scenarios while ensuring that knowledge of old categories is retained. For instance, voice assistants like Siri frequently encounter new event types, such as pandemics, need to understand and provide health-protecting information based on the user’s latest intent. Monaikul et al. (2021).
@@
Due to constraints imposed by storage limitations and privacy concerns, there exists a shortage of training data about the old categories He and Zhu (2022). Additionally, the manual relabeling of all categories within the new training dataset would incur substantial costs and time investment De Lange et al. (2021); Bang et al. (2021). Consequently, the model undergoes continuous updates using a freshly acquired dataset comprising the new categories. As depicted in Figure 1, the model undergoes training based on the current ground-truth label and undergoes testing using the full ground-truth label.
The incremental sequence labeling task faces a significant challenge known as the catastrophic forgetting problem, as extensively discussed in previous studies McCloskey and Cohen (1989); Robins (1995); Goodfellow et al. (2013); Kirkpatrick et al. (2017); Zheng et al. (2024). This issue manifests as semantic shifts, leading to a decrease in the discriminative power of entity classes Zhang et al. (2023b); Ma et al. (2023). In this paper, we decompose the problem into two primary semantic shifts in the incremental sequence task: E2O and O2E. The first semantic shift, E2O, arises from the presence of non-entities, potential old entities (mislabelled as non-entities), and new entities in the new dataset. Progress has been made in addressing E2O through methods falling into three categories: (1) Methods based on knowledge distillation: For instance, RDP proposes a knowledge distillation loss incorporating inter-task relations Zhang et al. (2023b). At the same time, CFPD introduces a pooled feature distillation loss to alleviate catastrophic forgetting Zhang et al. (2023a). (2) Methods based on pseudo-labels: OCILNER utilizes class prototypes to label new data Ma et al. (2023), and CPFD employs old models to label predictions of new data. (3) Methods based on freezing models: Examples include ICE Liu and Huang (2023), which freezes the backbone model and old classifiers to maintain the stability of the old classes at the expense of learning new classes.
@@ 2 Related Work
Incremental Learning The model continually acquires new tasks intending to achieve optimal performance on tasks previously learned Gepperth and Hammer (2016); Wu et al. (2019); van de Ven et al. (2022); Zheng et al. (2023b). There are three main categories of current incremental learning methods: regularization-based, rehearsal-based, and architecture-based. Regularization-based methods place constraints on model weights Kirkpatrick et al. (2017); Zenke et al. (2017), representations of intermediate layer features Hou et al. (2019); Douillard et al. (2020), and output probabilities Li and Hoiem (2017); Zheng et al. (2023a). Rehearsal-based methods overcome forgetting by saving some of the data containing the old classes for learning with the new classes Lopez-Paz and Ranzato (2017); Shin et al. (2017). Alternatively, architecture-based approaches involve dynamically expanding the network structure to allow for more data as new classes are added Hou et al. (2018); Yan et al. (2021).
@@
Methods for incremental sequence labeling tasks can be categorized into distillation-based, rehearsal-based, and other approaches. Distillation-based methods encompass ExtendNER Monaikul et al. (2021), which is the pioneer in applying knowledge distillation to incremental sequence labeling task, RDP Zhang et al. (2023b) with a relational distillation approach, and CPFD Zhang et al. (2023a) utilizing pooled features distillation loss. CFNER Zheng et al. (2022) introduces a causal framework for extracting new causal effects in entities and non-entities. Rehearsal-based approaches include KCN Cao et al. (2020) and KD+R+K Yu et al. (2021), both employing rehearsal samples to address the class imbalance and catastrophic forgetting in incremental event detection. L&R Xia et al. (2022) proposes a learn-and-review framework by training a new backbone model and a generative model simultaneously, generating synthetic samples of the old class to be trained with new samples. OCILNER Ma et al. (2023) uses rehearsal samples to compute class feature centers as class prototypes, generates an entity-oriented feature space through comparative learning, and annotates new data with pseudo-labels using class prototypes. Other methods encompass span-based and freezing model-based approaches, among others.
@@ 3 Problem Formulation
Formally, the objective of incremental sequence labeling is to acquire knowledge through a series of tasks $\mathcal{T}=\{\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{N}\}$. Each task contains its dataset $\mathcal{D}_{t}=\{(x^{i},y^{i})|y^{i}\in\mathcal{Y}_{t}\}$ where $(x^{i},y^{i})$ is a pair formed by the input token sentence and the label corresponding to each token in the sentence and $\mathcal{Y}_{t}$ stands for the current label set. Notably, $y^{i}$ only labels the token corresponding to the current task $t$, and the other tokens are labeled as O class (potential old entities $\mathcal{Y}_{1:t}$, and unseen entities $\mathcal{Y}_{t+1:N}$). At task $t$ ($t>1$), the new model $\mathcal{M}_{t}$ learns only from the new dataset and is expected to perform well on the learned classes $\bigcup_{i=1}^{t}\mathcal{Y}_{i}$.
@@ 4 Method
In this section, we systematically address the catastrophic forgetting problem by decomposing it into two distinct semantic shift challenges (Section 4.1). Subsequently, we present a comprehensive framework designed to address these semantic shifts individually, focusing on E2O in Section 4.2 and O2E in Section 4.3. The overarching goal is to effectively mitigate the catastrophic forgetting problem, as illustrated in Figure 4.
@@ 4.1 Two semantic shift problems
E2O refers to the model incorrectly categorizing entities as non-entities during the learning process. This misclassification stems from the incremental sequence labeling task, where only new entities are labeled in the new dataset, potentially causing old entities to be erroneously labeled as non-entities. For instance, in Figure 3, the name "Amy" is mistakenly labeled as a non-entity. This misclassification induces a gradual shift in the semantics of old entities towards non-entities, leading to a blurred boundary between the two classes. Several previous approaches have addressed this bias issue. Methods like RDP focus on designing improved distillation techniques to maintain the stability of the model’s old entities. Similarly, OCILNER utilizes comparative learning to obtain a more discriminative feature space, clarifying the classification boundaries between entities and non-entities. These strategies aim to mitigate the impact of E2O, ensuring a more accurate preservation of entity semantics during incremental sequence labeling tasks.
@@
O2E signifies the model incorrectly labeling non-entities or old entities as new entities during the learning process. As seen in Figure 3, our observations indicate that the model maintains good discrimination between old entities. However, Figure 2 shows a bias towards new entities in predictions during incremental learning. Our research identifies two key contributing factors to this bias.
@@ 4.2 Solving E2O problem via knowledge distillation
Through Eq.1, "Amy" in Figure 4 corrects the current model’s incorrect labeling via the output probabilities provided by the old model, thus maintaining discriminative properties between old entities.
@@ 4.3.1 Debiasing in Ordinary Cross Entropy
where $y_{i}$ denotes the label of the new entity for the current incremental step $t$. Figure 2 shows that confusion matrix of previous method at incremental step 4. It clearly shows that most predictions are biased towards the recent entity (class 4). We find that such a bias can be found in the cross-entropy loss function. When learning new entities, the model’s gradient update for old entities is defined as:
@@ 4.3.2 Learning with Prototypes
The class prototypes of old entities serve two essential purposes: firstly, they participate in optimization alongside new entities in each batch, ensuring a balanced optimization process among entities. Secondly, these class prototypes act as anchors in the feature space, mitigating the issue of over-labeling new entities. As depicted in Figure 4, the introduction of old prototypes reduces the potential over-labeling of new entities, enhancing the precision of new entity learning.
@@
Our approach differs from OCILNER’s approach, which uses prototypes in two ways:(1) OCILNER’s approach stores old samples for calculating prototypes. However, in this paper, we only use the training data in each incremental step for calculating prototypes and do not introduce replay samples. (2) OCILNER uses prototypes to label new datasets and adopts a cosine similarity as the threshold for entity labeling. However, in this paper, we found that some of the real non-entities also have a high cosine similarity with entities, which can easily produce wrong labeling for real non-entities and exacerbate semantic migration from entities to non-entities.
@@
Table 1: Comparisons with state-of-the-art methods on the i2b2 dataset using the bert-base-cased model. The best results are highlighted in bold and the second best results are underlined. The average of each incremental step is provided in Figure 5.
@@ 5.1 Experimental Setup
Settings We sort the above slices according to initial letter and train them in a FG-a-PG-b manner. FG means that the pre-trained model is trained with a entity types as the initial model, and PG means that the initial model is trained with b entity types at each following incremental step.
@@
Figure 6: The comparison between our method and previous state-of-the-art methods on nine incremental learning settings. We report the MacroF1 score after learning the final task. The detailed results are provided in Table 1, 8, 7, 9, and 6.
Implementation Details We use bert-base-cased model from HuggingFace Wolf et al. (2019) as backbone, with a hidden dimension of $d=768$. In addition to this, we have included supplemental experiments with the roberta-base model on the relevant datasets in Appendix E. We use the AdamW Loshchilov and Hutter (2018) optimizer, with learning rate $1e^{-6}$ and $1e^{-3}$ for backbone and classifier. We report the mean and standard deviation results over five runs.
@@ 5.2 Results and Analysis
Comparisons with State-Of-The-Art To validate the effectiveness of our approach, we conducted exhaustive experiments on the i2b2, OntoNotes5, and MAVEN datasets. We used the Finetune Only (FT) approach as a lower bound for comparison. Table 1, 8 displays the results of the experiments conducted on i2b2. Due to space limitations, we provide the results on MAVEN in Table 6 and OntoNotes5 in Table 7, 9. In detail, we show the experimental results under nine incremental learning settings through Figure 6. Our method consistently outperforms the previous state-of-the-art method in multiple settings, from FG-1-PG-1 to FG-8-PG-2. The poor performance of the previous method may be attributed to the ignorance of O2E.
@@
As shown in Figure 7, during the learning process, the previous method, ExtendNER, confuses new entities with non-entities due to O2E and old entities with non-entities due to E2O. Both of them together lead to poor prediction results of the model. We have effectively mitigated the above problems through our framework IS3, which strikes a good balance between maintaining old entities and learning new ones.
((a)) ExtendNER
((b)) IS3 (Ours)
To further demonstrate the effectiveness of our method, we visualize the feature representation through T-SNE Van der Maaten and Hinton (2008). As shown in Figure 9, the ExtendNER method faces serious E2O and O2E problems, with new entities and non-entities overwriting old ones, leading to catastrophic forgetting. Our method successfully addresses the issue of semantic bias that arises when the model learns a new task.
@@
Hyper-Parameter Analysis Figure 10 shows the results of different hyper-parameter choices on OntoNotes5 with the setting FG-1-PG-1. We consider two hyper-parameters: the correction factor in the debiased cross-entropy loss $\delta$ and the weight of the prototype loss $\beta$. The results show that $\delta$ around 0.5 reaches the best result, indicating that a moderate penalty effect reduction favors model performance. As $\beta$ keeps increasing, it makes the model overfit for the old prototype, leading to a decrease in model performance.
Case Study We provide an example in Figure 8 to demonstrate that the previous method suffers from an O2E offset when learning a new entity ORG, overwriting the old entity EVENT as a new entity. Simultaneously, the model inherits past O2E issues (labeling [O] as [DATE]). Additionally, it suffers from E2O, which fails to recognize the old entity accurately. Our method effectively balances these two offset problems and is more conducive to model learning.